# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v4

@@ B.2 Add Gaussian Noise to Raw Images
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&H(\mathcal{T};\bm{X})-H(\mathcal{T};\bm{X+\epsilon})\\ =&\frac{1}{2}\log|\bm{\Sigma_{Y|X}}|-\frac{1}{2}\log|\bm{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y|X}}|}{|\bm{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\bm{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (55)
Borrow the equations from the case of Gaussian noise added to the latent space, we have:
$\bigtriangleup S(\mathcal{T},\bm{\epsilon})=\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}$ (56)
Clearly, the introduction of Gaussian noise to each pixel in the original images has a detrimental impact on the task. Note that some studies have empirically shown that adding Gaussian noise to partial pixels of input images may be beneficial to the learning task Li (2022) Zhang et al. (2023).
@@ C.1 Inject Linear Transform Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\bm{Z})=&H(\mathcal{T};\bm{Z})-H(\mathcal{T};\bm{Z}+Q\bm{Z})\\ =&H(\bm{Y};\bm{Z})-H(\bm{Z})-(H(\bm{Y};\bm{Z}+Q\bm{Z})-H(\bm{Z}))\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}+Q\bm{Z})\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma}_{\bm{Z}}||\bm{\Sigma_{Y}}-\bm{\Sigma_{YZ}}{\bm{\Sigma_{Z}^{-1}}}\bm{\Sigma_{ZY}}|}{|\bm{\Sigma}_{(I+Q)\bm{Z}}||\bm{\Sigma_{Y}-\Sigma_{YZ}}\bm{\Sigma^{-1}_{Z}}\bm{\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (57)
@@
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\bm{Z})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (58)
@@
Rank Deficiency Cases. To avoid causing a rank deficiency of $I+Q$, we need to figure out the conditions that lead to rank deficiency. Here we show a simple case causing the rank deficiency. When the matrix $Q$ is a backward identity matrix Horn & R. (2012),
@@
$\mathrm{HermiteForm}(I+Q)_{i}=\bm{0},\quad\exists i\in\left[1,k\right]$ (62)
where index $i$ is the row index, in this paper, the row index starts from $1$, and $\mathrm{HermiteForm}$ is the Hermite normal form Kannan & Bachem (1979).
Full Rank Cases. Except for the rank deficiency cases, $I+Q$ has full rank and is invertible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows Shores (2007).
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\bm{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case of adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise into the embeddings. To this end, we demonstrate that linear transform can be positive noise.
@@ C.1.1 Optimal Quality Matrix of Linear Transform Noise
The determinant of the above equation is 0, but it breaks the first constraint of $\ rank(I+Q)=k$. However, by adding a small constant into the diagonal, and minus another constant by other elements, we can get:
@@
$Q_{optimal}=\mathrm{diag}\left(\frac{1}{k+1}-1,\dots,\frac{1}{k+1}-1\right)+\frac{1}{k+1}\bm{1}_{k\times k}$ (69)
@@
$\bigtriangleup S(\mathcal{T},Q\bm{X})_{upper}=(k-1)\log{(k+1)}$ (72)
@@ C.2 Add Linear Transform Noise to Raw Images
$\begin{split}H(\mathcal{T};\bm{X}+Q\bm{X})=&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X}+Q\bm{X})\log{p(\bm{Y}|\bm{X}+Q\bm{X})}\\ =&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|(I+Q)\bm{X})\log{p(\bm{Y}|(I+Q)\bm{X})}\end{split}$ (73)
where $I$ is an identity matrix, and $Q$ is derived from $I$ using elementary row operations. Assuming that the pixels of the raw images follow a Gaussian distribution. The conditional distribution of $\bm{Y}$ given $\bm{X}+Q\bm{X}$ is also multivariate subjected to the normal distribution, which can be formulated as:
$\bm{Y}|(I+Q)\bm{X}\sim\mathcal{N}(\mathbb{E}({\bm{Y}|(I+Q)\bm{X}}),var(\bm{Y}|(I+Q)\bm{X}))$ (74)
Since the linear transform matrix is invertible, applying the linear transform to $\bm{X}$ does not alter the distribution of the $\bm{X}$. It is straightforward to obtain:
$\bm{\mu}_{\bm{Y}|(I+Q)\bm{X}}=\bm{\mu_{Y}}+\bm{\Sigma_{YX}}\bm{{\Sigma}^{-1}_{X}}(I+Q)^{-1}((I+Q)X-(I+Q)\bm{\mu_{X}})$ (75)
$\bm{\Sigma}_{{(\bm{Y}|(I+Q)\bm{X}})}=\bm{\Sigma_{Y}-\Sigma_{YX}}\bm{\Sigma^{-1}_{X}}\bm{\Sigma_{XY}}$ (76)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},Q\bm{X})=&H(\mathcal{T};\bm{X})-H(\mathcal{T};\bm{X}+Q\bm{X})\\ =&\frac{1}{2}\log|\bm{\Sigma_{Y|X}}|-\frac{1}{2}\log|{\Sigma_{\bm{Y}|\bm{X}+Q\bm{X}}}|\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y|X}}|}{|{\Sigma_{\bm{Y}|\bm{X}+Q\bm{X}}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\bm{\Sigma_{Y}-\Sigma_{YX}}\bm{\Sigma^{-1}_{X}}\bm{\Sigma_{XY}}|}\\ =&0\end{split}$ (77)
@@ Appendix D Influence of Salt-and-pepper Noise on Task Entropy
Salt-and-pepper noise is a common type of noise that can occur in images due to various factors, such as signal transmission errors, faulty sensors, or other environmental factors Chan et al. (2005). Salt-and-pepper noise is often considered to be an independent process because it is a type of random noise that affects individual pixels in an image independently of each other Gonzales & Wintz (1987).
@@ D.1 Inject Salt-and-pepper Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\bm{Z})=&H(\mathcal{T};\bm{Z})-H(\mathcal{T};\bm{Z}\bm{\epsilon})\\ =&H(\bm{Y};\bm{Z})-H(\bm{Z})-(H(\bm{Y};\bm{Z}\bm{\epsilon})-H(\bm{Z}))\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}\bm{\epsilon})\\ =&-\sum_{\bm{Z}\in\mathcal{Z}}\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Z},\bm{Y})\log p(\bm{Z},\bm{Y})+\sum_{\bm{Z}\in\mathcal{Z}}\sum_{\bm{Y}\in\mathcal{Y}}\sum_{\bm{\epsilon}\in\mathcal{E}}p(\bm{Z\epsilon},\bm{Y})\log p(\bm{Z\epsilon},\bm{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{Z\epsilon},\bm{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{\epsilon})}\right]\\ =&-\mathbb{E}\left[\log\frac{1}{p(\bm{\epsilon})}\right]\\ =&-H(\bm{\epsilon})\end{split}$ (78)
@@ D.2 Add Salt-and-pepper Noise to Raw Images
$\begin{split}H(\mathcal{T};\bm{X}\bm{\epsilon})=&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X}\bm{\epsilon})\log{p(\bm{Y}|\bm{X}\bm{\epsilon})}\end{split}$ (79)
Since $\bm{\epsilon}$ is independent of $\bm{X}$ and $\bm{Y}$, the above equation can be expanded as:
$\begin{split}H(\mathcal{T};\bm{X}\bm{\epsilon})=&-\sum_{\bm{Y}\in\mathcal{Y}}\frac{p(\bm{Y},\bm{X}\bm{\epsilon})}{p(\bm{X})p(\bm{\epsilon})}\log{\frac{p(\bm{Y},\bm{X}\bm{\epsilon})}{p(\bm{X})p(\bm{\epsilon})}}\\ =&-\sum_{\bm{Y}\in\mathcal{Y}}\frac{p(\bm{Y},\bm{X})p(\bm{\epsilon})}{p(\bm{X})p(\bm{\epsilon})}\log{\frac{p(\bm{Y},\bm{X})p(\bm{\epsilon})}{p(\bm{X})p(\bm{\epsilon})}}\\ =&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X})\log{p(\bm{Y}|\bm{X})}\end{split}$ (80)
where
$\begin{split}p(\bm{X\epsilon},\bm{Y})=&p(\bm{X\epsilon}|\bm{Y})p(\bm{Y})\\ =&p(\bm{X}|\bm{Y})p(\bm{\epsilon}|\bm{Y})p(\bm{Y})\\ =&p(\bm{X}|\bm{Y})p(\bm{\epsilon})p(\bm{Y})\\ =&p(\bm{X},\bm{Y})p(\bm{\epsilon})\end{split}$ (81)
Therefore, the entropy change with salt-and-pepper noise is:
$\bigtriangleup S(\mathcal{T},Q\bm{X})=H(\mathcal{T};\bm{X})-H(\mathcal{T};\bm{X}\bm{\epsilon})=0$ (82)
Salt-and-pepper noise can not help reduce the complexity of the task, and therefore, it is considered a type of pure detrimental noise.

### v5

@@ B.2 Add Gaussian Noise to Raw Images
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H_{\mathcal{T}}(\boldsymbol{X})-H_{\mathcal{T}}(\boldsymbol{X+\epsilon})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X+\epsilon}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\end{split}$ (54)
Borrowing the equations from the case of Gaussian noise added to the latent space, we have:
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}$ (55)
Clearly, the introduction of Gaussian noise to each pixel in the original images has a detrimental impact on the task. Note that some studies have empirically shown that adding Gaussian noise to partial pixels of input images may be beneficial to the learning task [30] [69].
@@ C.1 Inject Linear Transform Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})=&H_{\mathcal{T}}(\boldsymbol{Z})-H_{\mathcal{T}}(\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y},\boldsymbol{Z}+Q\boldsymbol{Z})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Y},\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YZ}}{\boldsymbol{\Sigma_{Z}^{-1}}}\boldsymbol{\Sigma_{ZY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}}\boldsymbol{\Sigma^{-1}_{Z}}\boldsymbol{\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (56)
@@
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (57)
@@
Rank Deficiency Cases To avoid causing a rank deficiency of $I+Q$, we need to figure out the conditions that lead to rank deficiency. Here we show a simple case causing the rank deficiency. When the matrix $Q$ is a backward identity matrix [23],
@@
$\mathrm{HermiteForm}(I+Q)_{i}=\boldsymbol{0},\quad\exists i\in\left[1,k\right]$ (61)
where index $i$ is the row index, in this paper, the row index starts from $1$, and $\mathrm{HermiteForm}$ is the Hermite normal form [25].
Full Rank Cases Except for the rank deficiency cases, $I+Q$ has full rank and is invertible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows [52].
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})>0$ when Q=0.5IQ=-0.5I. The constraints are satisfied. This is just a simple case of adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise into the embeddings. To this end, we demonstrate that linear transform can be positive noise.
@@ C.1.1 Optimal Quality Matrix of Linear Transform Noise
The determinant of the above equation is 0, but it breaks the first constraint of rank(I+Q)=k\ rank(I+Q)=k. However, by adding a small constant into the diagonal, and minus another constant by other elements, we can get:
@@
$Q_{optimal}=\mathrm{diag}\left(\frac{1}{k+1}-1,\dots,\frac{1}{k+1}-1\right)+\frac{1}{k+1}\boldsymbol{1}_{k\times k}$ (68)
@@
$\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})_{upper}=(k-1)\log{(k+1)}$ (71)
@@ C.2 Add Linear Transform Noise to Raw Images
$\begin{split}H_{\mathcal{T}}(\boldsymbol{X}+Q\boldsymbol{X})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X})\log{p(\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X})}\\ =&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|(I+Q)\boldsymbol{X})\log{p(\boldsymbol{Y}|(I+Q)\boldsymbol{X})}\end{split}$ (72)
where $I$ is an identity matrix, and $Q$ is derived from $I$ using elementary row operations. Assuming that the pixels of the raw images follow a Gaussian distribution. The conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{X}+Q\boldsymbol{X}$ is also multivariate subjected to the normal distribution, which can be formulated as:
$\boldsymbol{Y}|(I+Q)\boldsymbol{X}\sim\mathcal{N}(\mathbb{E}({\boldsymbol{Y}|(I+Q)\boldsymbol{X}}),var(\boldsymbol{Y}|(I+Q)\boldsymbol{X}))$ (73)
Since the linear transform matrix is invertible, applying the linear transform to $\boldsymbol{X}$ does not alter the distribution of the $\boldsymbol{X}$. It is straightforward to obtain:
$\boldsymbol{\mu}_{\boldsymbol{Y}|(I+Q)\boldsymbol{X}}=\boldsymbol{\mu_{Y}}+\boldsymbol{\Sigma_{YX}}\boldsymbol{{\Sigma}^{-1}_{X}}(I+Q)^{-1}((I+Q)X-(I+Q)\boldsymbol{\mu_{X}})$ (74)
$\boldsymbol{\Sigma}_{{(\boldsymbol{Y}|(I+Q)\boldsymbol{X}})}=\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}$ (75)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})=&H_{\mathcal{T}}(\boldsymbol{X})-H_{\mathcal{T}}(\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X}}|-\frac{1}{2}\log|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|{\Sigma_{\boldsymbol{Y}|\boldsymbol{X}+Q\boldsymbol{X}}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&0\end{split}$ (76)
@@ Appendix D Influence of Salt-and-pepper Noise on Task Entropy
Salt-and-pepper noise is a common type of noise that can occur in images due to various factors, such as signal transmission errors, faulty sensors, or other environmental factors [8]. Salt-and-pepper noise is often considered to be an independent process because it is a type of random noise that affects individual pixels in an image independently of each other [20].
@@ D.1 Inject Salt-and-pepper Noise in Latent Space
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=$ $H_{\mathcal{T}}(\boldsymbol{Z})-H_{\mathcal{T}}(\boldsymbol{Z}\boldsymbol{\epsilon})$
$=$ $H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Y},\boldsymbol{Z\epsilon})$
$=$ $H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y},\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z}))$
$=$ $\mathbb{E}\!\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\!\left[\log\frac{1}{p(\boldsymbol{Z\epsilon},\boldsymbol{Y})}\right]$
$\leq$ $\mathbb{E}\!\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\!\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]\;=\;0,$
@@ D.2 Add Salt-and-pepper Noise to Raw Images
$\begin{split}H_{\mathcal{T}}(\boldsymbol{X}\boldsymbol{\epsilon})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X}\boldsymbol{\epsilon})\log{p(\boldsymbol{Y}|\boldsymbol{X}\boldsymbol{\epsilon})}\end{split}$ (77)
Since $\boldsymbol{\epsilon}$ is independent of $\boldsymbol{X}$ and $\boldsymbol{Y}$, and $I(\boldsymbol{Y};\boldsymbol{X\epsilon})\leq I(\boldsymbol{Y};\boldsymbol{X})\ \Longrightarrow\ H(\boldsymbol{Y}\mid\boldsymbol{X})\leq H(\boldsymbol{Y}\mid\boldsymbol{X\epsilon})$. The entropy change satisfies:
$\Delta S(\mathcal{T},\boldsymbol{\epsilon})=H(\boldsymbol{Y}\mid\boldsymbol{X})-H(\boldsymbol{Y}\mid\boldsymbol{X\epsilon})\leq 0$ (78)
Therefore, salt-and-pepper noise can not help reduce the complexity of the task, and therefore, it is considered a type of pure detrimental noise.
@@
Figure 3: Differnet Noise Injection Modes. The green tick mark means noise injection operations are implemented, while the red cross means no noise injection. We adopt the Mode 1 in this work. Modes 2 and 3 are for future exploration.